in the united states 2/5 of them start with M,A,or N how can you use equivelent fractions to find out how many states this is?
1
Expert's answer
2013-02-04T09:22:19-0500
There are 50 states totally. So, there are

50 * 2/5 = 100/5 = 20

states that start with M, A, or N.